OpenStudy (anonymous):

Neophyte Need help! Geometry Find the Midpoint of Line CA C=(-2,0) A=(2,4) With the work please(:

OpenStudy (anonymous):

add each component and divide by 2

OpenStudy (anonymous):

-2+2=0 0/2=0

OpenStudy (anonymous):

(0, )

OpenStudy (anonymous):

0+4=4 4/2=2 (0,2)

OpenStudy (anonymous):

Thank you.